OpenStudy (anonymous):

Can someone see if i'm doing this right ? The formula : F=P(1+r/n)^nt Odessa deposited $1,347.80 into a savings account with an interest rate of 3.3% compounded monthly. About how long will it take for the account to be worth $5,000? F=1,347.80 P=5,000 N=12 R=3.3% 0.033 1,347.80(1+0.033/12)^12t ?

OpenStudy (anonymous):

I menat 1,347.80=5,000(1+0.033/12)^12t

OpenStudy (anonymous):

You were on the right track in that last line in your problem statement. You would set that equal to 5000 and then take logs.

OpenStudy (anonymous):

In your last post, you need to switch the positioning of those 2 dollar amounts. This is because 5000 is what you end up with, not what you start out with.

OpenStudy (anonymous):

Is this starting to make sense now, @NaCl ?

OpenStudy (anonymous):

no i'm a visual leaner .

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

5000=1347.80(1+0.033/12)^12t ?

OpenStudy (anonymous):

Yes! You've got it now!

OpenStudy (anonymous):

An easy way to think about the 2 dollar amounts is to see what happens to the deposit of "one dollar". You would put "1" where you have 1347.80 and see what it grows to.

OpenStudy (anonymous):

would it be 39 years, 10 months ?

OpenStudy (anonymous):

In that wolfram equation, you have to put 12t in parentheses because exponentiation takes precedence over multiplication. However, you did get the right answer with: 39.780788780031041623654225812091 Good job!

OpenStudy (anonymous):

Good luck to you in all of your studies! @NaCl

OpenStudy (anonymous):

Sonya currently has an account balance of $1,533.43. Her initial deposit on the account was $962 and it earned 3.3% simple interest. How long has Sonya held the account? 30 years 11 years 48 years 18 years

OpenStudy (phi):

simple interest is Interest = Principal * rate * time to do your problem, figure out how much interest Sonya received (total - start) then solve for time in the above equation

OpenStudy (anonymous):

@phi I did I=962*3.3%*48=1523.808 ? is 48 years right ?

OpenStudy (phi):

that would be the amount of simple interest you get after 48 years the amount in the account would be the interest plus how much you start with. Step 1: if you start with 962, and after some years have 1,533.43 how much interest did you get ?

OpenStudy (anonymous):

3.3 %

OpenStudy (phi):

that is a rate. Think like this: you start with 1 apple, and some time later you have 3 apples. How many apples were added ?

OpenStudy (phi):

Step 1: if you start with 962, and after some years have 1,533.43 how many dollars were added to the account?

OpenStudy (anonymous):

o.O idk

OpenStudy (anonymous):

Your end account balance is 1533.43 where you started with 962, giving you 571.43 total interest because 1533.43 - 962 = 571.43 That means 962 dollars generated 571.43 dollars in interest. Each dollar generated: 571.43 / 962 = 0.594002079 dollars in interest (0.033) times (t years) = 0.594002079 t = 0.594002079 / 0.033 = 18 years

OpenStudy (phi):

are you saying you can't figure out how much bigger 1533.43 is than 962?

OpenStudy (phi):

as tcarroll shows, you get 571.43 in interest because 962 + 571.43= 1533.43 now you can use the formula Interest= Principal * rate * time 571.43 = 962* 3.3% * t to solve for t

OpenStudy (anonymous):

This explanation of @phi s is really good. It's basically the same as mine. I added a step which may or may not be necessary that has to do with what one dollar generates in interest. I tend to do that because it is rather fundamental to consider one dollar.

OpenStudy (anonymous):

Anyway, @NaCl you can look at my post of 2 previous to see how I arrived at 18 years.

OpenStudy (anonymous):

Thank you both for you time @tcarroll010 @phi

OpenStudy (anonymous):

You are very welcome! Anytime! If you are ok now with these 2 problems, then close it when you are ready, and if you have any new problems, just start a brand-new thread. Otherwise, if you are still in need of any further explanation on this problem, we can still kick it around.